# DiffNMR2: NMR Guided Sampling Acquisition Through Diffusion Model Uncertainty

### v1

@@ Abstract
Diffusion Model, Nuclear Magnetic Resonance, Uncertainty
Figure 1: a) Regular NMR experiment. A 2D NMR resonance map in the time-time domain $(t_{e},t_{a})$ is produced by recording the response of nuclei subjected to two-pulse sequences, where each row is obtained with a different time delay (called evolution time: $t_{e}$) between the two pulses and $t_{a}$ represent the acquisition time. This map is first transformed into a frequency-time representation via a 1D Fourier Transform along the acquisition time axis: $(t_{e},t_{a})\to(t_{e},f_{a})$. A second 1D Fourier Transform along the evolution time axis produces the final spectrum: $(t_{e},f_{a})\to(f_{e}f_{a})$. b) Non-uniform sampling (NUS). Differing from the regular NMR experiment, Non-Uniform Sampling selects a subset of the evolution times (Row-wise Scan). After 1D Fourier Transform on the acquisition time axis, existing methods such as Compressed Sensing (Kazimierczuk & Orekhov, 2011) and Low-Rank Approximation (Qu et al., 2015) are applied to reconstruct the frequency-time domain (Row Inpainting). c) Guided sampling workflow. Differing from NUS, we propose a novel dynamic sampling strategy by leveraging the diffusion model uncertainty. This is the first attempt using the diffusion model during NMR acquisition. The novelties are highlighted in blue texts.
@@ 1 Introduction
Nuclear Magnetic Resonance (NMR) spectroscopy is a powerful tool for identifying the molecular structures, providing insights into the arrangement of atoms within complex molecules. It has become an indispensable tool in chemistry, biochemistry, and structural biology from its introduction in 1938 (Giunta & Mainz, 2020). However, acquiring NMR spectra requires long acquisition times to reconstruct high-resolution information (Fig. 1). Traditional acquisition methods can be prohibitively slow in applications involving large biomolecules or when time is bounded, such as in drug discovery (Pellecchia et al., 2002, 2008) or metabolomics (Markley et al., 2017; Emwas et al., 2019). For example, acquiring a 4D NMR spectrum (CCNOESY) of the phycobilisome linker polypeptide domain of CpcC takes up to 90 hours (Ramelot et al., 2011).
Non-uniform sampling (NUS) addresses this challenge by randomly sub-sampling the scanned signal while enabling high-quality spectral reconstructions (Fig. 1). Instead of sampling every data point in a regular pattern, NUS collects a subset of data points based on a pseudo-random or optimized sampling schedule (Marvasti, 1993). The advantage of NUS lies in its capacity to significantly accelerate acquisition, but its effectiveness relies on robust reconstruction algorithms capable of handling incomplete data. Several methods have been developed for reconstructing spectra from NUS data, including Low-Rank Approximation (LR) (Qu et al., 2015), Compressed Sensing (CS) (Kazimierczuk & Orekhov, 2011), and Deep Learning (DL)-based methods (Qu et al., 2020; Zhan et al., 2024; Zheng et al., 2022; Karunanithy & Hansen, 2021).
In this study, we introduce an AI-based approach that dynamically adapts the acquisition strategy. By integrating acquisition and reconstruction, the method leverages the uncertainty of the diffusion model (i.e., the degree of variability in the model’s reconstruction) to refine sampling in real-time, significantly reducing the acquisition burden. Using a real-life protein 2D NMR dataset (Klukowski et al., 2024), we demonstrate that our method outperforms state-of-the-art sampling techniques and reconstruction methods, delivering accurate high-resolution spectral reconstructions from substantially reduced datasets. This work establishes a new paradigm in NMR spectroscopy by combining the strengths of NUS and advanced machine learning, enabling rapid, high-quality spectral analysis for time-sensitive applications.
@@ 2.3 Non-Uniform Sampling (NUS)
Non-Uniform Sampling (NUS) is a technique used to improve the efficiency of 2D NMR data acquisition by selectively sampling a subset of the evolution times ($t_{e}$) rather than collecting the full Nyquist grid. Instead of acquiring data at every increment of $t_{e}$, only a fraction of the points are sampled at random. One standard strategy is the Sine-Weighted Poisson-Gap sampling (Hyberts et al., 2010), which assumes a Poisson distribution for the gap size between evolution times.
The reconstruction of the full spectrum from sparsely sampled data requires advanced computational algorithms. One prominent method is Low-Rank Approximation (LRA) (Qu et al., 2015), which represents the data as a Hankel matrix and approximates it with a low-rank reconstruction that best fits the sampled points. While computationally efficient, the LRA struggles to capture highly complex spectral features or systems with noise, leading to artifacts in the reconstruction. Compressed Sensing (Kazimierczuk & Orekhov, 2011) is another standard that leverages the sparsity of NMR signals in the time-time domain to reconstruct the complete dataset. However, CS requires careful tuning of regularization parameters and may exhibit limitations in reconstructing spectra with dense peaks or overlapping signals.
In recent literature, Qu et al. (Qu et al., 2020) introduced a dense Convolutional Neural Network (CNN) (LeCun et al., 2015) with a spectrum consistency block for Poisson-Gap NUS data, achieving promising results for 2D and 3D spectra. Zhan et al. (Zhan et al., 2024) employed SEPSNet, an attention-based network, to enhance pure shift NMR reconstruction from a 2D resonance map. Zheng et al. (Zheng et al., 2022) developed PS-ResNet, a ResNet-based model, for denoising 1D spectra. Karunanithy et al. (Karunanithy & Hansen, 2021) applied dilated convolutions through Fidnet, inspired by WaveNet, for synthetic NMR data analysis. Despite these advancements, these deep learning methods share a common limitation: they are primarily trained on synthetic datasets, which may not fully capture the complexity and variability of real-life NMR data. As a result, their performance has not consistently surpassed the established CS baseline for NUS reconstruction, particularly when applied to experimental datasets (Yan et al., 2024).
@@ 3.1 NUS Reconstruction with Diffusion Model inpainting
Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), are a class of neural networks originally inspired by non-equilibrium thermodynamics processes (Sohl-Dickstein et al., 2015). These models are trained by progressively destroying data with noise and then learning to reverse the process and regenerate the original data. The first phase (i.e., the diffusion process) consists of generating noisy data such that:
@@
The second phase (i.e., denoising process), involves training a UNet (Ronneberger et al., 2015) to minimize the difference between the predicted noise and the actual noise added in the forward process. This can be formulated as optimizing the Mean Reconstruction Error loss:
@@
The same model and architecture can also be used for inpainting in the context of incomplete image reconstruction. In our experiments, we perform the inpainting using a diffusion model trained for denoising combined with the Repaint pipeline (Lugmayr et al., 2022). This pipeline works by combining the information of the unmasked area during the denoising step. This is illustrated in Fig. 2.
@@ 3.3 Data
Recent advancements in diffusion models for image generation rely on massive datasets containing hundreds of millions of images (Schuhmann et al., 2021; Rombach et al., 2022). In contrast, large-scale datasets for 2D NMR protein spectra are scarce, with only a few public repositories available. The standard approach has been to train models on synthetic spectra (Zhan et al., 2024; Qu et al., 2020; Karunanithy & Hansen, 2021; Zheng et al., 2022), which often fail to capture the complex noise patterns and signal variations present in real-life data. This limitation also explains why Compressed Sensing (Kazimierczuk & Orekhov, 2011) and Low-Rank Approximation (Qu et al., 2015) remain the state-of-the-art methods for reconstructing real-world NMR spectra. To address this gap, our study focuses on training models exclusively on real-life NMR spectra, providing a more realistic and reliable basis for evaluation.
In this study, we use the 100-protein NMR spectra dataset (Klukowski et al., 2024), comprising 1329 spectra in 2D, 3D, and 4D formats. These spectra, derived from 100 proteins, were sampled using NMR machines operating at frequencies ranging from 600 to 950 MHz. We augmented the dataset by including 2D spectra obtained by projecting higher-dimensional (3D or 4D) spectra into 2D representations. This technique, commonly used in NMR workflows for visualization purposes, increases the total number of samples to over 3500. The process maintains the integrity of the original signals while significantly enriching the dataset for training. Finally, the spectra are transformed with a 1D Inverse Fourier Transform to obtain their representation in the frequency-time domain where each row corresponds to a 1D NMR spectrum, and to an evolution time.
@@ 4 Experiments and Results
After inpainting the frequency-time representation, we obtain the reconstructed spectrum with a 1D Fourier Transform on the evolution time axis. All experiments and benchmark metrics are based on the quality of this final reconstructed spectrum since this is the output analyzed by NMR users. We simulate the partial acquisition by masking rows of the frequency-time representation, which is a fair approximation applied in the partial NMR acquisition (Qu et al., 2020; Karunanithy & Hansen, 2021; Zheng et al., 2022; Yan et al., 2024; Zhan et al., 2024).
@@ 4.1.1 Global Metrics
Given our goal of developing a method to reliably reconstruct the raw signal outputted by an NMR machine before any expert signal processing. We use the Mean Squared Error (MSE, lower is better) and the coefficient of determination (R2, higher is better) to globally evaluate the quality of the predictions. MSE is a standard reconstruction metric (Hyberts et al., 2010) that quantifies the magnitude of prediction errors, ensuring accuracy in capturing critical spectral features like peak intensities and positions. R² complements this by assessing how well the model explains the variance in the data, providing a normalized measure of fit. Together, these metrics ensure a robust evaluation of the method’s precision and reliability.
@@ 4.3 Baselines
We compare our versions with two other sampling methods from the literature, Uniform random sampling, which samples evolution time from a uniform distribution, and Poisson-Gap sampling (Hyberts et al., 2010). This last method has been established as a state-of-the-art sampling method regarding reconstruction quality. It operates by sampling gaps between evolution time from a Poisson-Gap distribution weighted with a sinus function that gives more importance to lower frequencies.
@@ 5 Discussion
In this section, we analyze the spectra reconstructed from the independent test set, focusing on two key aspects: the comparison of sampling strategies (i.e., Guided Sampling, Poisson-Gap Sampling, Uniform Sampling) and the comparison of inpainting methods (i.e., Diffusion model, Compressed Sensing (Kazimierczuk & Orekhov, 2011), Low-Rank Approximation (Qu et al., 2015)). Together, these comparisons provide a comprehensive understanding of the strengths of our approach in improving NMR spectrum spectroscopy.
@@ 5.1 Cross-sampling and cross-model comparisons
Cross-model comparison. While CS (Kazimierczuk & Orekhov, 2011) and LR (Qu et al., 2015) use Non-Uniform Sampling, their reconstruction relies on specialized algorithms that fall short of the diffusion model (Guided Sampling, Poisson-Gap, and Uniform Sampling) in terms of MSE, R2 and Hallucination Ratio. LR reconstructs fewer Missed Peaks than the diffusion model after 15% of acquisition completion. As aforementioned, this is due to noise being misidentified as peaks.
@@ 5.3 Case study
Applied to the CCNOESY NMR spectrum of the phycobilisome linker polypeptide domain of CpcC - 2L8V - nuclei H1 H2 (Ramelot et al., 2011), the reconstructed results (displayed in Fig. 6) show the practical effects of the method choice. After collecting 10% of the pulse sequences, Uniform sampling, CS and LR show a high proportion of noise, which severely compromise the peak analysis. Poisson-Gap strongly improves the signal reconstruction and retains most of the signal information, but contains a small amount of missed peak and hallucinations. The Guided Sampling approach (especially GS400) contains the least hallucinations and recovers most of the signal.
@@ 5.4 Wall Time Comparison
Using the average wall times for fully sampled 2D NMR acquisitions (nmr, ), we estimated the total wall time (diffusion model inference + pulse sequence acquisition) required for each NUS strategy to achieve a hallucination ratio below 10% (Fig. 8).

### v2

@@ Abstract
Keywords Diffusion Model, Nuclear Magnetic Resonance, Uncertainty
Figure 1: a) Regular NMR experiment. A 2D NMR resonance map in the time-time domain $(t_{e},t_{a})$ is produced by recording the response of nuclei subjected to two-pulse sequences, where each row is obtained with a different time delay (called evolution time: $t_{e}$) between the two pulses and $t_{a}$ represent the acquisition time. This map is first transformed into a frequency-time representation via a 1D Fourier Transform along the acquisition time axis: $(t_{e},t_{a})\to(t_{e},f_{a})$. A second 1D Fourier Transform along the evolution time axis produces the final spectrum: $(t_{e},f_{a})\to(f_{e}f_{a})$. b) Non-uniform sampling (NUS). Differing from the regular NMR experiment, Non-Uniform Sampling selects a subset of the evolution times (Row-wise Scan). After 1D Fourier Transform on the acquisition time axis, existing methods such as Compressed Sensing [1] and Low-Rank Approximation [2] are applied to reconstruct the frequency-time domain (Row Inpainting). c) Guided sampling workflow. Differing from NUS, we propose a novel dynamic sampling strategy by leveraging the diffusion model uncertainty. This is the first attempt using the diffusion model during NMR acquisition. The novelties are highlighted in blue texts.
@@ 1 Introduction
Nuclear Magnetic Resonance (NMR) spectroscopy is a powerful tool for identifying the molecular structures, providing insights into the arrangement of atoms within complex molecules. It has become an indispensable tool in chemistry, biochemistry, and structural biology from its introduction in 1938 [3]. However, acquiring NMR spectra requires long acquisition times to reconstruct high-resolution information (Fig. 1). Traditional acquisition methods can be prohibitively slow in applications involving large biomolecules or when time is bounded, such as in drug discovery [4, 5] or metabolomics [6, 7]. For example, acquiring a 4D NMR spectrum (CCNOESY) of the phycobilisome linker polypeptide domain of CpcC takes up to 90 hours [8].
Non-uniform sampling (NUS) addresses this challenge by randomly sub-sampling the scanned signal while enabling high-quality spectral reconstructions (Fig. 1). Instead of sampling every data point in a regular pattern, NUS collects a subset of data points based on a pseudo-random or optimized sampling schedule [9]. The advantage of NUS lies in its capacity to significantly accelerate acquisition, but its effectiveness relies on robust reconstruction algorithms capable of handling incomplete data. Several methods have been developed for reconstructing spectra from NUS data, including Low-Rank Approximation (LR) [2], Compressed Sensing (CS) [1], and Deep Learning (DL)-based methods [10, 11, 12, 13].
In this study, we introduce an AI-based approach that dynamically adapts the acquisition strategy. By integrating acquisition and reconstruction, the method leverages the uncertainty of the diffusion model (i.e., the degree of variability in the model’s reconstruction) to refine sampling in real-time, significantly reducing the acquisition burden. Using a real-life protein 2D NMR dataset [14], we demonstrate that our method outperforms state-of-the-art sampling techniques and reconstruction methods, delivering accurate high-resolution spectral reconstructions from substantially reduced datasets. This work establishes a new paradigm in NMR spectroscopy by combining the strengths of NUS and advanced machine learning, enabling rapid, high-quality spectral analysis for time-sensitive applications.
@@ 2.3 Non-Uniform Sampling (NUS)
Non-Uniform Sampling (NUS) is a technique used to improve the efficiency of 2D NMR data acquisition by selectively sampling a subset of the evolution times ($t_{e}$) rather than collecting the full Nyquist grid. Instead of acquiring data at every increment of $t_{e}$, only a fraction of the points are sampled at random. One standard strategy is the Sine-Weighted Poisson-Gap sampling [15], which assumes a Poisson distribution for the gap size between evolution times.
The reconstruction of the full spectrum from sparsely sampled data requires advanced computational algorithms. One prominent method is Low-Rank Approximation (LRA) [2], which represents the data as a Hankel matrix and approximates it with a low-rank reconstruction that best fits the sampled points. While computationally efficient, the LRA struggles to capture highly complex spectral features or systems with noise, leading to artifacts in the reconstruction. Compressed Sensing [1] is another standard that leverages the sparsity of NMR signals in the time-time domain to reconstruct the complete dataset. However, CS requires careful tuning of regularization parameters and may exhibit limitations in reconstructing spectra with dense peaks or overlapping signals.
In recent literature, Qu et al. [10] introduced a dense Convolutional Neural Network (CNN) [16] with a spectrum consistency block for Poisson-Gap NUS data, achieving promising results for 2D and 3D spectra. Zhan et al. [11] employed SEPSNet, an attention-based network, to enhance pure shift NMR reconstruction from a 2D resonance map. Zheng et al. [12] developed PS-ResNet, a ResNet-based model, for denoising 1D spectra. Karunanithy et al. [13] applied dilated convolutions through Fidnet, inspired by WaveNet, for synthetic NMR data analysis. Despite these advancements, these deep learning methods share a common limitation: they are primarily trained on synthetic datasets, which may not fully capture the complexity and variability of real-life NMR data. As a result, their performance has not consistently surpassed the established CS baseline for NUS reconstruction, particularly when applied to experimental datasets [17].
@@ 3.1 NUS Reconstruction with Diffusion Model inpainting
Denoising Diffusion Probabilistic Models (DDPM) [18], are a class of neural networks originally inspired by non-equilibrium thermodynamics processes [19]. These models are trained by progressively destroying data with noise and then learning to reverse the process and regenerate the original data. The first phase (i.e., the diffusion process) consists of generating noisy data such that:
@@
The second phase (i.e., denoising process), involves training a UNet [20] to minimize the difference between the predicted noise and the actual noise added in the forward process. This can be formulated as optimizing the Mean Reconstruction Error loss:
@@
The same model and architecture can also be used for inpainting in the context of incomplete image reconstruction. In our experiments, we perform the inpainting using a diffusion model trained for denoising combined with the Repaint pipeline [21]. This pipeline works by combining the information of the unmasked area during the denoising step. This is illustrated in Fig. 2.
@@ 3.3 Data
Recent advancements in diffusion models for image generation rely on massive datasets containing hundreds of millions of images [22, 23]. In contrast, large-scale datasets for 2D NMR protein spectra are scarce, with only a few public repositories available. The standard approach has been to train models on synthetic spectra [11, 10, 13, 12], which often fail to capture the complex noise patterns and signal variations present in real-life data. This limitation also explains why Compressed Sensing [1] and Low-Rank Approximation [2] remain the state-of-the-art methods for reconstructing real-world NMR spectra. To address this gap, our study focuses on training models exclusively on real-life NMR spectra, providing a more realistic and reliable basis for evaluation.
In this study, we use the 100-protein NMR spectra dataset [14], comprising 1329 spectra in 2D, 3D, and 4D formats. These spectra, derived from 100 proteins, were sampled using NMR machines operating at frequencies ranging from 600 to 950 MHz. We augmented the dataset by including 2D spectra obtained by projecting higher-dimensional (3D or 4D) spectra into 2D representations. This technique, commonly used in NMR workflows for visualization purposes, increases the total number of samples to over 3500. The process maintains the integrity of the original signals while significantly enriching the dataset for training. Finally, the spectra are transformed with a 1D Inverse Fourier Transform to obtain their representation in the frequency-time domain where each row corresponds to a 1D NMR spectrum, and to an evolution time.
@@ 4 Experiments and Results
After inpainting the frequency-time representation, we obtain the reconstructed spectrum with a 1D Fourier Transform on the evolution time axis. All experiments and benchmark metrics are based on the quality of this final reconstructed spectrum since this is the output analyzed by NMR users. We simulate the partial acquisition by masking rows of the frequency-time representation, which is a fair approximation applied in the partial NMR acquisition [10, 13, 12, 17, 11].
@@ 4.1.1 Global Metrics
Given our goal of developing a method to reliably reconstruct the raw signal outputted by an NMR machine before any expert signal processing. We use the Mean Squared Error (MSE, lower is better) and the coefficient of determination (R2, higher is better) to globally evaluate the quality of the predictions. MSE is a standard reconstruction metric [15] that quantifies the magnitude of prediction errors, ensuring accuracy in capturing critical spectral features like peak intensities and positions. R² complements this by assessing how well the model explains the variance in the data, providing a normalized measure of fit. Together, these metrics ensure a robust evaluation of the method’s precision and reliability.
@@ 4.3 Baselines
We compare our versions with two other sampling methods from the literature, Uniform random sampling, which samples evolution time from a uniform distribution, and Poisson-Gap sampling [15]. This last method has been established as a state-of-the-art sampling method regarding reconstruction quality. It operates by sampling gaps between evolution time from a Poisson-Gap distribution weighted with a sinus function that gives more importance to lower frequencies.
@@ 5 Discussion
In this section, we analyze the spectra reconstructed from the independent test set, focusing on two key aspects: the comparison of sampling strategies (i.e., Guided Sampling, Poisson-Gap Sampling, Uniform Sampling) and the comparison of inpainting methods (i.e., Diffusion model, Compressed Sensing [1], Low-Rank Approximation [2]). Together, these comparisons provide a comprehensive understanding of the strengths of our approach in improving NMR spectrum spectroscopy.
@@ 5.1 Cross-sampling and cross-model comparisons
Cross-model comparison. While CS [1] and LR [2] use Non-Uniform Sampling, their reconstruction relies on specialized algorithms that fall short of the diffusion model (Guided Sampling, Poisson-Gap, and Uniform Sampling) in terms of MSE, R2 and Hallucination Ratio. LR reconstructs fewer Missed Peaks than the diffusion model after 15% of acquisition completion. As aforementioned, this is due to noise being misidentified as peaks.
@@ 5.3 Case study
Applied to the CCNOESY NMR spectrum of the phycobilisome linker polypeptide domain of CpcC - 2L8V - nuclei H1 H2 [8], the reconstructed results (displayed in Fig. 6) show the practical effects of the method choice. After collecting 10% of the pulse sequences, Uniform sampling, CS and LR show a high proportion of noise, which severely compromise the peak analysis. Poisson-Gap strongly improves the signal reconstruction and retains most of the signal information, but contains a small amount of missed peak and hallucinations. The Guided Sampling approach (especially GS400) contains the least hallucinations and recovers most of the signal.
@@ 5.4 Wall Time Comparison
Using the average wall times for fully sampled 2D NMR acquisitions [24], we estimated the total wall time (diffusion model inference + pulse sequence acquisition) required for each NUS strategy to achieve a hallucination ratio below 10% (Fig. 8).